# ACEGEN: Reinforcement learning of generative chemical agents for drug discovery

### v2

@@ 1 Introduction
Reinforcement learning (RL) has emerged as a possible solution to explore this chemical space 4, 5 with an increasing focus on efficiency 6, 7. RL 8 consists of a family of machine learning algorithms that use feedback as a learning signal. Through a trial and error process, RL algorithms can adapt their search strategy to the specific characteristics of compounds associated with desirable properties, and guide the generative models to propose novel molecules with similar characteristics. These methods use a customized reward function that assigns values to molecules, tailored specifically for the application or research question being addressed. To navigate the molecular space and find the most highly rewarded molecules, RL generative agents alternate between two phases: a data collection phase and a learning phase. During the data collection, molecules are generated by the generative model and scored based on the reward function. In the agent learning phase, the collected data is utilized to enhance the agent’s ability to generate interesting and relevant compounds. By leveraging this feedback loop, the generative model can be continually improved to generate compounds with increasingly desirable properties. This is particularly interesting when the available chemical space cannot be enumerated and filtered using the scoring function, either because the scoring functions are very slow or because the search space is very large.
@@
Currently, available implementations for drug discovery using reinforcement learning (RL) rely heavily on custom code 9, 10. This approach tends to favor redundancy, complexity, and limited efficiency, making it difficult to incorporate a diverse set of solutions. However, the RL community has already developed solutions to address these challenges. TorchRL, a comprehensive RL library 11, offers well-tested, independent state-of-the-art RL components. In this work, we adopt TorchRL’s components as building blocks to assemble efficient and reliable drug discovery agents—a practice already successfully applied in diverse domains such as drone control 12 and combinatorial optimization 13. This approach also fosters algorithmic research by enabling the encapsulation of new ideas within new components, which can seamlessly integrate with existing ones. TorchRL operates within the PyTorch ecosystem 14, ensuring not only high-quality standards but also maintenance over time and future development.
@@
To showcase the advantages of ACEGEN, we implement and evaluate well-known language-based algorithms for drug design. Language models can learn complex patterns in text and generate novel sequences, including molecular structures 15, 16. Studies have focused on utilizing the Simplified Molecular Input Line Entry System (SMILES) 17, a string notation system used to represent chemical structures in a compact and standardized manner. Moreover, several prior studies have demonstrated the benefits of combining generative language models and RL 4, 18, 7. Therefore, we re-implement three REINFORCE-based algorithms: REINFORCE 19, REINVENT 4, and AHC 6. Additionally, we adapt general RL algorithms such as A2C 20 and PPO 21 to our problem setting. The REINVENT and AHC methods utilize some level of experience replay; therefore, we also incorporate experience replay into REINFORCE. Similarly, we test PPOD 22, a PPO-based algorithm adapted for experience replay.
@@ 2.1 Reinforcement learning setting
Reinforcement learning tasks are formalized as Markov Decision Processes (MDPs) 8. Formally, an MDP is described by a quintuple, denoted as $\langle S,A,R,P,\rho_{0}\rangle$. Here, $S$ represents the set of all possible states in the problem space, $A$ denotes the set of valid actions available to the agent, and $R:S\times A\times S\to\mathbb{R}$ defines the reward function, which assigns numerical values to transitions from one state to another based on the action taken. The function $P:S\times A\to\mathcal{P}(S)$ indicates the transition probability, where $P(s_{t+1}|s,a)$ represents the probability of transitioning to state $s_{t+1}$ from the current state $s$ under action $a$. Lastly, $\rho_{0}$ signifies the initial state distribution.
This approach naturally extends to the problem of sequentially designing molecules. In this context, a parameterized reinforcement learning (RL) policy, denoted as $\pi_{\theta}$, guides molecule design by successively processing the current state representing a partially built molecule and selecting actions $a_{t}$ to make changes to the molecule’s structure at each step until reaching a terminal state. The reward function evaluates the quality of the molecule, assigning a scalar value either at each generative step or at the end of the process when a complete compound is produced. Within this framework, $\pi_{\theta}(a_{t}|s_{t})$ denotes the probability of the policy function, parameterized by $\theta$, taking action $a_{t}$ in state $s_{t}$. The sequence of actions needed to construct a molecule, often referred to as an episode, is represented by $\tau$. $P(\tau|\theta)$ indicates the probability of a trajectory $\tau$ given the policy parameters $\theta$. Furthermore, $R(\tau)$ represents the cumulative sum of rewards over the trajectory $\tau$.
The goal of RL algorithms is to optimize the parameters of the policy $\pi_{\theta}$ to maximize $R(\tau)$. Different methods within the family of policy gradient methods 8, such as REINFORCE 19, A2C (Advantage Actor-Critic) 20, and PPO (Proximal Policy Optimization) 21, are commonly used for this task.
Additionally, various techniques can aid in training. Reward shaping involves modifying the reward function to provide additional feedback to the agent during training, potentially speeding up learning by providing more informative signals. Experience replay involves storing past experiences in a buffer and randomly sampling from it during training, which can improve sample efficiency and stabilize learning. Moreover, incorporating penalty terms, such as a Kullback-Leibler (KL) divergence loss term, can be utilized to encourage the agent to stay close to a reference or prior policy, helping to maintain stability and control the exploration-exploitation trade-off during training. Finally, ranking and selecting only the best K molecules in each given batch can improve efficiency. ACEGEN integrates all the aforementioned algorithms and additional techniques to form a comprehensive suite of state-of-the-art methods for molecular generation.
@@ 2.2 Chemical language generative models
One possible approach to sequentially designing molecules is through string representations. Molecular string representations 23 convert molecular graphs into strings and vice versa, allowing molecules to be encoded using specific syntactical rules. This approach formulates the task of molecular generation as a natural language processing (NLP) 24 problem, and the policy models used to sequentially generate molecules represented as strings are called Chemical Language Models (CLMs) 25. In this context, the resulting action space is discrete, with each action denoted as a token. These actions are integrated by concatenating them to form a joint string representation, producing the next state. Each episode begins with a single special start token and can last for a varying number of steps. The episode ends when the agent chooses another special token called the stop token.
ACEGEN currently provides an environment for language model experimentation including SMILES 17, DeepSMILES 26, SELFIES 27, AtomInSmiles 28 and SAFE 29 grammars, complemented by a user-friendly vocabulary class.  Appendix A A in the appendix illustrates how the vocabulary and environment can be easily created and utilized for data generation in ACEGEN.
In practice, CLMs are first trained unsupervised on a bulk of unlabeled data, aimed at learning to generate valid molecules, also referred to as pre-training, which is achieved through the application of the teacher enforcing method. 30. The pre-trained CLM is the starting policy that can be further trained to optimize a specific objective with RL, similar to what happens in large language models. Moreover, this prior policy can be used as an anchor point from which the new RL policy should not deviate excessively.
ACEGEN currently provides pre-trained models for several architectures. More specifically, we provide GRU 31 policies pre-trained on two datasets: ChEMBL 32 and ZINC 16. We also provide weights for an LSTM 33 policy pre-trained on ChEMBL 32 and for a GPT2 34 pre-trained on the REAL Database REAL lead-like compounds 35.
While our repository currently provides methods to create LSTM 33, GRU 31, GPT2 34 , Llama2 36 and Mamba 37 policies, practitioners can similarly include any other architectures of their choice taking advantage of current and future developments in NLP. Users can pre-train their policies with external code and integrate them without any need to modify the internal scripts of the library. Our repository contains a step-by-step tutorial that explains how to do the integration.
Finally, ACEGEN also provides its own pre-training script to train language models, either from ACEGEN or custom, to generate valid molecules through the teacher enforcing method 30. Our code is engineered to adapt to the available computational resources, whether a single GPU or a distributed setup spanning multiple machines and GPUs. This adaptability allows us to efficiently train on datasets of large size.
@@ 2.3 Scoring and Evaluation of Molecules
Scoring functions must reflect real-world drug design scenarios and be flexible enough to apply to a range of drug design challenges. Often, overly simplistic objective functions are used for optimization, such as penalized logP 38, or complex solutions tailored to one particular generative model 9. ACEGEN allows the integration of custom scoring functions by providing the flexibility to define them as Python methods that accept strings and return numerical values. Alternatively, ACEGEN is optionally compatible with the MolScore 39 library, which offers a broad range of drug design relevant scoring functions, diversity filters, support for curriculum learning and also contains a benchmarking mode including MolOpt 10, GuacaMol 3, and others.  Appendix B B in the appendix showcases how both scoring options can be accommodated within our codebase. Additionally, the library offers a detailed tutorial, guiding users through the process of implementing and incorporating custom scoring functions seamlessly into the entire workflow to train agents with them.
@@ 2.4 RL Agents Training
ACEGEN provides training for RL agents utilizing the following methods: REINFORCE 19, REINVENT 9, AHC 6, A2C 20, and PPO 21, as well as an adapted version of the PPOD algorithm 22. We deviate from the exact implementation of PPOD by omitting the use of an initial expert demonstration, replaying only episodes with high rewards (instead of episodes with high-value predictions), and employing a fixed amount of replay data per batch. All methods are fully configurable, allowing the use of custom scoring functions and models beyond those already provided by the library.
REINVENT and AHC can be considered extensions of the REINFORCE algorithm. Specifically, REINVENT incorporates reward shaping, experience replay, and a penalty for high-likelihood sequences into the basic REINFORCE framework. Additionally, AHC ranks and selects the top K molecules in each batch of data. We find that using the same reward shaping is not compatible with advantage-based methods, and the high-likelihood penalty term is detrimental. Therefore, for A2C, PPO, and PPOD, we introduce a term to the loss function based on the Kullback-Leibler divergence (KL) between the actor policy and a prior policy. This term serves to penalize the policy for deviating too much from a prior, similar in concept to reward shaping in REINVENT and AHC. Incorporating KL constraints is a common practice in research papers aligning language models with custom reward functions, as seen in examples utilizing human feedback 40, 41, 42.
@@ 2.5 De-novo, decorative and fragment-linking generation
In drug discovery pipelines, generating molecules from scratch may not always suffice. To meet diverse requirements, ACEGEN offers multiple sampling modes. PromptSMILES 43 is a simple method enabling constrained molecule generation using models pre-trained solely with teacher enforcement on full SMILES, making it possible to generate molecules while adhering to specific chemical sub-structures. Specifically, in addition to de-novo generation, ACEGEN scripts allow for easy configuration of scaffold decoration and fragment-linking molecule generation modes. Constrained sampling modifies only the data collection behavior, operating independently of all other agent components, in line with the TorchRL philosophy. Tutorials on performing constrained generation are provided within the repository.
@@ 3.1 Benchmarking RL performance
To assess different RL algorithms, we have compared performance on the Practical Molecular Optimization (MolOpt) benchmark 10 as implemented in MolScore 39. This benchmark encompasses 23 distinct tasks associated with different objectives to be optimized within a budget of 10,000 molecules. All algorithms use the same GRU policy model architecture as in the original MolOpt benchmark, implemented in ACEGEN and pretrained on a curated subset of ChEMBL 32. We compared REINFORCE 19, REINVENT with the hyperparameters from the original paper 9, REINVENT-MolOpt with optimized hyperparameters for the MolOpt benchmark 10, and AHC with the hyperparameters from the original paper 6. Note that we utilize experience replay in all of the aforementioned algorithms Additionally, we also test A2C 20, PPO 21 which does not utilize any experience replay, and PPOD 22 which is an adaptation of PPO to accommodate experience replay. All algorithms are ACEGEN implementations, and the hyperparameter values used for all of them are provided in our repository. As a sanity check, we compared our implementation of REINVENT with the implementation used in the MolOpt paper. Our simplified re-implementation achieved better results with faster training. These results are shown in Appendix C C in the appendix.
@@
As proposed by the MolOpt benchmark authors, Table Table 1 shows algorithm performance on the AUC of the top 10 molecules as an indication of the sample efficiency of identifying 10 desirable molecules with respect to the objective, a representative figure that might be carried forward to later stages of drug design. This shows that PPOD is state-of-the-art with respect to sample efficiency, followed by REINVENT-MolOpt and then PPO. To achieve a higher level understanding of performance, Table 2 shows a selection of metrics aimed to measure maximum performance, efficiency, and exploration (a common trade-off with exploitation and hence efficiency in RL). These results show that REINVENT-MolOpt achieves maximum performance as measured by the average top 10 molecules, PPOD maximum efficiency as measured by AUC of the top 10 molecules, and AHC maximum exploration as measured by the number of unique compounds generated.
Table 2: Algorithm comparison on a selection of metrics analyzing the performance on the MolOpt benchmark. Valid is a sanity check that measures the proportion of valid molecules generated. Top-10 AUC is a measure of sample efficiency as shown in 1. Top-10 Avg is a measure of the absolute best 10 molecules achieved with the budget. Unique is a proxy for exploration which measures the proportion of unique molecules generated. Basic and target chemistry filters (B&T-CF) are the proportion of molecules after filtering out highly idiosyncratic molecules in general or with respect to the pretraining dataset. The B&T-CF Top-10 AUC and B&T-CF Top-10 Avg are recalculated on this subset while also enforcing that the 10 molecules identified offer diverse solutions (Div). Lastly, B&T-CF Diversity is a proxy measure of exploration via sphere exclusion diversity 44 on this subset. Each algorithm was run 5 times with different seeds, and results were averaged and then summed over the 23 tasks. A perfect score for all metrics is 23.
@@
Valid 21.77 21.78 21.74 21.45 18.93 21.78 21.52
Top-10 Avg 15.85 15.69 17.43 16.09 15.67 15.63 17.10
Top-10 AUC 13.67 13.60 15.65 13.91 14.27 14.65 15.80
Unique 22.28 22.63 13.68 22.68 18.38 9.47 10.21
B&T-CF 14.34 14.74 7.00 13.77 7.82 5.81 5.71
B&T-CF Top-10 Avg (Div) 14.83 14.79 16.06 15.25 14.54 14.60 15.78
B&T-CF Top-10 AUC (Div) 12.85 12.81 14.61 13.11 13.32 13.67 14.60
B&T-CF Diversity (SEDiv@1k) 17.39 18.19 10.10 17.55 14.24 8.47 7.96

### v3

@@ 1 Introduction
Reinforcement learning (RL) has emerged as a possible solution to explore this chemical space 4, 5 with an increasing focus on efficiency 6, 7. RL 8 is a family of machine learning algorithms that use feedback as a learning signal to guide a decision-making process. Thus, RL algorithms can adapt a molecule building decision making progress to achieve certain characteristics of molecules resulting in novel molecules with desirable properties. This effectively constitutes a search strategy of chemical space. RL requires a reward function that assigns a value to a molecule, tailored specifically for the application or research question being addressed. The RL algorithm then seeks to maximize this value, therefore, adapting the navigation through the decision making progress and hence chemical space. This is particularly interesting when the available chemical space cannot be enumerated and filtered using the scoring function, either because the scoring functions are very slow or because the search space is very large.
There are some key differences between the application of RL classically and the application to search chemical space for drug design. Classically RL algorithms are developed and evaluated in game theory which usually has a well defined environment and a clear objective, for example, complete this level, or win the game. This is a huge discrepancy compared to drug discovery and design 9, where the objective is multi-faceted (an efficacious, non-toxic, and bioavailable molecule) and measured by proxies that poorly define the true objective e.g., optimizing the predicted binding affinity to a protein compared to the true objective of sufficiently perturbing a cellular pathway. This is unfortunately necessary due to the cost of running experimental assays. In light of this expected error between optimized proxy and true objective, it is important to propose diverse solutions 10 to increase the chance of identifying a molecule of interest with all desirable properties for further development - another contrast to classical RL where a single successful solution is often sufficient. Despite these challenges, RL has shown promising preliminary results with successful application to drug design across a number of different molecular representations and model architectures 11.
@@
However, currently available implementations for drug discovery using reinforcement learning (RL) rely heavily on custom code 12, 13. This approach tends to favor redundancy, complexity, and limited efficiency, making it difficult to incorporate a diverse set of solutions. However, the RL community has already developed solutions to address these challenges. TorchRL 14, a comprehensive RL library, offers well-tested, independent state-of-the-art RL components. In this work, we adopt TorchRL’s components as building blocks to assemble efficient and reliable drug discovery agents—a practice already successfully applied in diverse domains such as drone control 15 and combinatorial optimization 16. This approach also fosters algorithmic research by enabling the encapsulation of new ideas within new components, which can seamlessly integrate with existing ones. TorchRL operates within the PyTorch ecosystem 17, ensuring not only high-quality standards but also maintenance over time and future development.
@@
To showcase the advantages of ACEGEN, we implement and evaluate well-known language-based algorithms for drug design. An overview of the workflow for ACEGEN implementation is shown in Figure 1. Language models can learn complex patterns in text and generate novel sequences, including molecular structures 18, 19. Studies have focused on utilizing the Simplified Molecular Input Line Entry System (SMILES) 20, a string notation system used to represent chemical structures in a compact and standardized manner. Moreover, several prior studies have demonstrated the benefits of combining generative language models and RL 4, 21, 7. Therefore, we re-implement three REINFORCE-based algorithms: REINFORCE 22, REINVENT 4, and AHC 6. Additionally, we adapt general RL algorithms such as A2C 23 and PPO 24 to our problem setting. The REINVENT and AHC methods utilize some level of experience replay; therefore, we also incorporate experience replay into REINFORCE. Similarly, we test PPOD 25, a PPO-based algorithm adapted for experience replay.
Figure 1: General overview of any of the ACEGEN implementations. Different ACEGEN implementations vary in the algorithms used, and allow to customize the generative models and the scoring functions.
@@ 2.1 Reinforcement learning setting
Reinforcement learning (RL) tasks are formalized as Markov Decision Processes (MDPs) 8, described by the quintuple $\langle S,A,R,P,\rho_{0}\rangle$. Here, $S$ is the set of all possible states in the problem space, $A$ is the set of valid actions available to the agent, and $R:S\times A\times S\to\mathbb{R}$ is the reward function, which assigns numerical value to the transition from one state to the next given the action taken. The function $P:S\times A\to\mathcal{P}(S)$ is the transition probability, where $P(s_{t+1}|s,a)$ is the probability of transitioning to state $s_{t+1}$ from the current state $s$ under action $a$. Lastly, $\rho_{0}$ signifies the initial state distribution.
This can be applied to the problem of designing molecules sequentially. In this context, a parameterized RL policy $\pi_{\theta}$ can navigate molecule design by selecting actions (molecular edits) $a_{t}$ at a given state $s_{t}$ (partially built molecule) until molecule building is terminated. The reward function assigns a scalar value that captures desirability, either at each step or upon termination of molecule building. Where $\pi_{\theta}(a_{t}|s_{t})$ denotes the probability of taking action $a_{t}$ in state $s_{t}$ from the policy function parameterized by $\theta$. Meanwhile, $\tau$ represents the episode, or sequence of actions needed to construct a molecule. Thus, $P(\tau|\theta)$ is the probability of the full trajectory $\tau$ given the policy parameters $\theta$, and $R(\tau)$ is the cumulative sum of rewards over the trajectory $\tau$.
The goal of policy-based RL algorithms is to optimize the parameters of the policy $\pi_{\theta}$ to maximize $R(\tau)$. Different methods within the family of policy gradient methods 8, such as REINFORCE 22, A2C (Advantage Actor-Critic) 23, and PPO (Proximal Policy Optimization) 24, are commonly used for this task.
Additionally, various techniques can aid in training. Reward shaping modifies the reward function to provide additional feedback to the agent during training, potentially providing more informative signals. Experience replay stores past experiences in a buffer and randomly samples from it during training, which can improve sample efficiency and stabilize learning. Incorporating penalty terms, such as a Kullback-Leibler (KL) divergence loss term, can encourage the agent to stay close to a reference or prior policy, helping to maintain stability. Finally, ranking and selecting only the best K molecules in each given batch can improve efficiency. ACEGEN integrates all the aforementioned algorithms and additional techniques to form a comprehensive suite of state-of-the-art methods for molecular generation.
@@ 2.2 Chemical language generative models
Molecular string representations 26 convert molecular graphs into strings and vice versa. This representation therefore formulates the task of molecular generation as a natural language processing (NLP) 27 problem, and the policy models used to sequentially generate molecules represented as strings are called Chemical Language Models (CLMs) 28. In this context, the resulting action space is discrete, with each action $a$ denoted as a token, and each non-terminal state $s_{t}$ a partially complete SMILES string. Each episode begins with a single special start token, e.g., "GO" and can last for a varying number of steps. The episode ends when the agent chooses another special token called the stop token e.g., "EOS".
ACEGEN currently provides an environment for language model experimentation including SMILES 20, DeepSMILES 29, SELFIES 30, AtomInSmiles 31 and SAFE 32 grammars, complemented by a user-friendly vocabulary class. Section A in the supporting information illustrates how the vocabulary and environment can be easily created and utilized for data generation in ACEGEN.
CLMs are first trained unsupervised on a bulk of unlabeled data, aimed at learning to generate valid molecules, achieved through the application of the teacher enforcing method. 33. The pre-trained CLM is the starting policy function $\pi_{\theta}$ that can be further trained to optimize a specific objective with RL. Moreover, this prior policy can be used as an anchor point from which the new RL policy should not deviate excessively.
ACEGEN currently provides pre-trained models for several architectures. We provide GRU 34 policies pre-trained on two datasets: ChEMBL 35 and ZINC 19. We also provide an LSTM 36 policy pre-trained on ChEMBL 35, and a GPT2 37 policy pre-trained on the Enamine Database REAL lead-like compounds 38. Our repository provides ready-to-use architectures for LSTM 36, GRU 34, GPT2 37 , Llama2 39 and Mamba 40 policies. Moreover, users can include any other architectures of their choice following our step-by-step tutorial that explain how to do the integration, which is possible without modifying any ACEGEN internals.
Finally, ACEGEN also provides pre-training script to train language models, either from ACEGEN or custom. Our code is engineered to adapt to the available computational resources, whether a single GPU or a distributed setup spanning multiple machines and GPUs. This adaptability allows to efficiently train models on datasets of large size.
@@ 2.3 Scoring and Evaluation of Molecules
Scoring functions must reflect real-world drug design scenarios and be flexible enough to apply to a range of drug design challenges. Often, overly simplistic objective functions are used for optimization, such as penalized logP 41, or complex solutions tailored to one particular generative model 12. ACEGEN allows the integration of custom scoring functions by providing the flexibility to define them as Python methods that accept strings and return numerical values.
Section B in the supporting information showcases how scoring functions can be implemented easily in ACEGEN. Additionally, the library offers a detailed tutorial, guiding users through the process of implementing and incorporating custom scoring functions seamlessly into the entire workflow to train agents with them.
In this paper for benchmarking we have used the MolScore 42 library, which offers a broad range of drug design relevant scoring functions, diversity filters, support for curriculum learning and also contains a benchmarking mode including MolOpt 13, GuacaMol 3, and others.
@@ 2.4 RL Agents Training
ACEGEN provides training for RL agents utilizing the following methods: REINFORCE 22, REINVENT 12, AHC 6, A2C 23, and PPO 24, as well as an adapted version of the PPOD algorithm 25. We deviate from the exact implementation of PPOD by omitting the use of an initial expert demonstration, replaying only episodes with high rewards (instead of episodes with high-value predictions), and employing a fixed amount of replay data per batch. All methods are fully configurable, allowing the use of custom scoring functions and models beyond those already provided by the library.
REINVENT and AHC can be considered extensions of the REINFORCE algorithm. Specifically, REINVENT incorporates reward shaping, experience replay, and a penalty for high-likelihood sequences into the basic REINFORCE framework. Additionally, AHC ranks and selects the top K molecules in each batch of data. We find that using the same reward shaping is not compatible with advantage-based methods, and the high-likelihood penalty term is detrimental. Therefore, for A2C, PPO, and PPOD, we introduce a term to the loss function based on the Kullback-Leibler divergence (KL) between the actor policy and a prior policy. This term serves to penalize the policy for deviating too much from a prior, similar in concept to reward shaping in REINVENT and AHC. Incorporating KL constraints is a common practice in research papers aligning language models with custom reward functions, as seen in examples utilizing human feedback 43, 44, 45.
@@ 2.5 De-novo, decorative and fragment-linking generation
In drug discovery pipelines, generating molecules from scratch may not always suffice. To meet diverse requirements, ACEGEN offers multiple sampling modes. PromptSMILES 46 is a simple method enabling constrained molecule generation using models pre-trained solely with teacher enforcement on full SMILES, making it possible to generate molecules while adhering to specific chemical sub-structures. Specifically, in addition to de-novo generation, ACEGEN scripts allow for easy configuration of scaffold decoration and fragment-linking molecule generation modes. Constrained sampling modifies only the data collection behavior, operating independently of all other agent components, in line with the TorchRL philosophy. Tutorials on performing constrained generation are provided within the repository.
@@ 3.1 Benchmarking RL performance
To assess different RL algorithms, we have compared performance on the Practical Molecular Optimization (MolOpt) benchmark 13 as implemented in MolScore 42. This benchmark encompasses 23 distinct tasks associated with different objectives to be optimized within a budget of 10,000 molecules. All algorithms use the same GRU policy model architecture as in the original MolOpt benchmark, implemented in ACEGEN and pretrained on a curated subset of ChEMBL 35. We compared REINFORCE 22, REINVENT with the hyperparameters from the original paper 12, REINVENT-MolOpt with optimized hyperparameters for the MolOpt benchmark 13, and AHC with the hyperparameters from the original paper 6. Note that we utilize experience replay in all of the aforementioned algorithms Additionally, we also test A2C 23, PPO 24 which does not utilize any experience replay, and PPOD 25 which is an adaptation of PPO to accommodate experience replay. All algorithms are ACEGEN implementations, and the hyperparameter values used for all of them are provided in our repository. As a sanity check, we compared our implementation of REINVENT with the implementation used in the MolOpt paper. Our simplified re-implementation achieved better results with faster training. These results are shown in section C in the supporting information.
@@
As proposed by the MolOpt benchmark authors, Table Table 1 shows algorithm performance on the AUC of the top 10 molecules as an indication of the sample efficiency of identifying 10 desirable molecules with respect to the objective, a representative figure that might be carried forward to later stages of drug design. This shows that PPOD is state-of-the-art with respect to sample efficiency, followed by REINVENT-MolOpt and then PPO. To achieve a higher level understanding of performance, Table 2 and Figure 2a shows a selection of metrics aimed to measure maximum performance, efficiency, and exploration (a common trade-off with exploitation and hence efficiency in RL). These results show that REINVENT-MolOpt achieves maximum performance as measured by the average top 10 molecules, PPOD maximum efficiency as measured by AUC of the top 10 molecules, and AHC maximum exploration as measured by the number of unique compounds generated.
Table 2: Algorithm comparison on a selection of metrics analyzing the performance on the MolOpt benchmark. Valid is a sanity check that measures the proportion of valid molecules generated. Top-10 AUC is a measure of sample efficiency as shown in 1. Top-10 Avg is a measure of the absolute best 10 molecules achieved with the budget. Unique is a proxy for exploration which measures the proportion of unique molecules generated. Basic and target chemistry filters (B&T-CF) are the proportion of molecules after filtering out highly idiosyncratic molecules in general or with respect to the pretraining dataset. The B&T-CF Top-10 AUC and B&T-CF Top-10 Avg are recalculated on this subset while also enforcing that the 10 molecules identified offer diverse solutions (Div). Lastly, B&T-CF Diversity is a proxy measure of exploration via sphere exclusion diversity 47 on this subset. Each algorithm was run 5 times with different seeds, the average score and variance was then summed over the 23 tasks to report the final score and standard deviation of the summed variance. A perfect score for all metrics is 23.
@@
Valid 21.77 $\pm$ 0.03 21.78 $\pm$ 0.02 21.74 $\pm$ 0.05 21.45 $\pm$ 0.02 18.93 $\pm$ 0.51 21.78 $\pm$ 0.09 21.52 $\pm$ 0.15
Top-10 Avg 15.85 $\pm$ 0.10 15.69 $\pm$ 0.11 17.43 $\pm$ 0.17 16.09 $\pm$ 0.10 15.67 $\pm$ 0.22 15.63 $\pm$ 0.27 17.10 $\pm$ 0.19
Top-10 AUC 13.67 $\pm$ 0.07 13.60 $\pm$ 0.08 15.65 $\pm$ 0.14 13.91 $\pm$ 0.08 14.27 $\pm$ 0.14 14.65 $\pm$ 0.23 15.80 $\pm$ 0.14
Unique 22.28 $\pm$ 0.10 22.63 $\pm$ 0.06 13.68 $\pm$ 0.31 22.68 $\pm$ 0.07 18.38 $\pm$ 0.50 9.47 $\pm$ 0.33 10.21 $\pm$ 0.35
B&T-CF 14.34 $\pm$ 0.18 14.74 $\pm$ 0.15 7.00 $\pm$ 0.27 13.77 $\pm$ 0.12 7.82 $\pm$ 0.37 5.81 $\pm$ 0.26 5.71 $\pm$ 0.22
B&T-CF Top-10 Avg (Div) 14.95 $\pm$ 0.12 14.83 $\pm$ 0.13 16.06 $\pm$ 0.16 15.25 $\pm$ 0.10 14.54 $\pm$ 0.18 14.60 $\pm$ 0.23 15.78 $\pm$ 0.15
B&T-CF Top-10 AUC (Div) 12.91 $\pm$ 0.06 12.85 $\pm$ 0.08 14.61 $\pm$ 0.13 13.11 $\pm$ 0.08 13.32 $\pm$ 0.12 13.67 $\pm$ 0.17 14.60 $\pm$ 0.14
B&T-CF Diversity (SEDiv@1k) 17.39 $\pm$ 0.15 18.19 $\pm$ 0.12 10.10 $\pm$ 0.31 17.55 $\pm$ 0.10 14.07 $\pm$ 0.47 7.61 $\pm$ 0.49 7.71 $\pm$ 0.39
Figure 2: Comparison of RL algorithms by radar plot visualization of metric performance for (a) the MolOpt benchmark reported in Table 2, for (b) the MolOpt benchmark with chemistry requirements explicitly in the reward signal as reported in Table 3, (c) the 5-HT2A case study as reported in Table 5, and (d) the REINFORCE ablation study as reported in Table 4. The legend at the top of the figure applies to sub-plots (a), (b), and (c). Subplot (d) has the legend beside.